# Iterative Inversion of Ill-Conditioned MIMO Channels Using Symmetric Rank-$1$ Regularization

### v3

@@ III-B Optimality Study
This result indicates the best matrix condition that can be possibly achieved via the SR-$1$R approach.
@@
IV Power Iteration-Assisted Approach for The Implementation of SR-$1$R
@@ IV-A The Proposed PIA Approach
Algorithm 1 Pseudo code of the PIA for the SR-$1$R
@@
6:  Form the matrix $\mathbf{R}$ via (1) and compute $\mathbf{R}^{-1}$ through the Schulz method specified in (II-B2);
@@
Finally, the proposed PIA approach for SR-$1$R implementation is summarized in Algorithm 1.
@@ IV-B Convergence of The PIA Approach
7:  Let $\lambda_{\textsc{old}}=\lambda$ and goto step $3$;
@@ IV-C Enhanced PIA Approach
[…]
$\widetilde{\lambda}(i)=\mbox{\boldmath$\mathcal{\phi}$}^{T}(i)\mathbf{c}^{\prime}.$ (77)
@@
[…]
@@ IV-D Scalability Analysis for SR-$1$R
In this subsection, we analyze the scalability of iterative linear precoding with the SR-$1$R in terms of both the overall complexity and the algorithm-depth. Then, we compare the SR-$1$R to the preconditioning baselines. The complexity analysis reflects the computational cost of the SR-$1$R; while the algorithm-depth reflects the minimum computing steps (i.e., the shortest processing time) when employing parallel computing [43]. This measure is particularly relevant because compatibility with parallel computing is a key advantage of iterative matrix inversion methods [1]. It will be shown that both the PIA/e-PIA algorithms exhibit favorable compatibility with parallel computing.
@@
  • Step $6$. This step belongs to the iterative matrix inversion. A good regularization/preconditioning technique accelerates this process. Each iteration costs two matrix-matrix multiplications and a matrix-matrix deduction. When the iteration nmber is $i$, this yields the complexity of $(2N^{3}+N^{2})i$ and the algorithm-depth of $2i\log_{2}N$. In Section V, we will demonstrate the computational latency when the iterative matrix inversion converges (see Table I).
@@
Summing up the aforementioned steps yields the overall complexity of the iterative linear precoding with SR-$1$R:
$(2i+1)N^{3}+(\tau_{0}+\tau_{N-1}+i+M+7)N^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(M+\log_{2}N+4)N.$ (81)
Since our primary goal is to handle ill-conditioned MIMO systems, it is reasonable to assume $N\approx M$. Moreover, both $\tau_{0}$ and $\tau_{N-1}$ are small constants in our implementation (see Section V-A). Therefore, the complexity level of the iterative linear precoding with the SR-$1$R is $\mathcal{O}(iN^{3})$.
Similarly, the overall algorithm-depth of the iterative linear precoding with the SR-$1$R is given by
$(2i+\tau_{0}+\tau_{N-1}+7)\log_{2}N+\log_{2}M,$ (82)
where the algorithm-depth level is $\mathcal{O}(i\log_{2}N)$. Notably, both the complexity and algorithm-depth are dominated by the iterative matrix inversion. This means using the SR-$1$R does not significantly increase the computational cost or latency of iterative linear precoding.
When employing the e-PIA, Algorithm 1 is performed in parallel for $L$ times. Therefore, the complexity level of the e-PIA is $\mathcal{O}(iLN^{3})$; while the algorithm-depth level remains the same as the PIA.
To compare the complexity and the algorithm-depth of the SR-$1$R to the preconditioning baselines, we focus on the parts where they are different. To this end, we exclude the iterative matrix inversion (Step 6 of Algorithm 1) and the additional steps of iterative linear precoding, as they are common to both the SR-$1$R and the preconditioning baselines. In this case, the complexity level of the SR-$1$R is $\mathcal{O}(N^{2})$, and the algorithm-depth level is $\mathcal{O}(\log_{2}N)$. In comparison, the GS and SSOR preconditioning require the inversion of a triangular matrix, i.e., a complexity of $\mathcal{O}(N^{2})$ [34]. This is usually done by the Gauss-Jordan elimination, and constitutes $N$ iterations [44]. Hence, the algorithm-depth is $\mathcal{O}(N)$.
In Section V, we will demonstrate the scalability of the SR-$1$R in terms of the computation latency when parallel computing is supported (see Fig. 5) as well as the overall latency of iterative linear precoding when RZF performance is achieved (see Table I).
@@ V Simulation Results and Discussion
In this section, computer simulations are carried out to demonstrate the performance of both the PIA and e-PIA approaches for the SR-$1$R. As an overview, the results are divided into five experiments (see Section V-C):
  • Experiment 1 investigates the performance of the proposed PIA/e-PIA approaches in comparison to the theoretical expectation in Theorem 1;
  • Experiment 2 investigates the scalability of the SR-$1$R in terms of computation latency;
  • Experiment 3 investigates the condition number improvement of the PIA/e-PIA approaches;
  • Experiment 4 compares the RZF to the ZF when using iterative linear precoding;
  • Experiment 5 investigates the iteration number of the PIA/e-PIA to achieve the average symbol error rate (SER) of RZF.
Although the SR-$1$R achieves good performance improvement as will be demonstrated in Section V-C), it is limited by the performance bound in (17). Therefore, as suggested by the reviewer, we show an additional example to demonstrate the potential of small-rank regularization in Experiment 6.
V-A Implementation of the SR-$1$R
@@
(83) means the increment approaches zero in two cases: when $i\to\infty$, or when $(\lambda_{n})/(\lambda_{0})\approx 1$. When $\mathbf{u}$ is badly initialized, i.e., $\lambda_{n}\approx\lambda_{0}$, it is hard to determine the convergence. This brings troubles to the PIA and e-PIA approaches to calculate $\hat{\mathbf{u}}_{N-1}$
@@ V-B Baselines and Simulation Setup
We consider three types of baselines as follows. Their names and abbreviations are highlighted as follows.
  • Iterative Matrix Inversion
    We use the Schulz method (see (II-B2)) [2].
  • Preconditioning
    1. 1.
      Jacobi preconditioning (see (14)) [23, Eq. (10)].
    2. 2.
      Gauss-Seidel (GS) preconditioning (see (15)) [6, Eq. (14)].
    3. 3.
      Symmetric successive over-relaxation (SSOR) (see (16)) [16, Section II.B].
  • Linear System Inverse
    1. 1.
      Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [32, Eq. (2.6)].
    2. 2.
      Modified randomized iterative detection algorithm (MRIDA) [9, Algorithm 2].
The iterative matrix inversion alone and the preconditioning are included to show the improvement of the SR-$1$R on the condition number (see Experiment 4$\sim$5). In the study of iterative linear precoding, the linear system inverse is also included since it serves the same purpose as iterative matrix inversion (see Experiment 5). Both the BFGS algorithm and the MRIDA have demonstrated remarkable reliability advantage in the recent studies [6, 9, 7, 8].
We consider two types of ELAA channel and two types of stationary ill-conditioned channel in the simulation. The channel model in [15] is adopted for the ELAA channel in appreciation of its comprehensive consideration of mixed LoS/non-LoS states, shadowing effects and path-loss. In ELAA systems, we consider the height of users and the base station to be $1.5$ m and $10$ m, respectively.
  • LoS-dominated ELAA: The ill condition arises from the correlation of the LoS paths. We consider $M=64$, $N=16$ (single user), and the vertical distance from the user to the ELAA is $40$ m.
  • Symmetric Mixed LoS/non-LoS ELAA: The ill condition arises from both the correlation of the LoS paths and the symmetric MIMO size. The MIMO size is configured as: $M=N=128$ ($8$ users with $16$ antennas each), and the vertical distance from users to the ELAA is $40$ m. The users are equally spaced along a parallel line to the base station antenna array with a maximal inter-user distance of $10$ m.
Apart from the ELAA channels, we also consider stationary channels where $M=N=128$ or $M=N=512$.
  • Symmetric i.i.d. Rayleigh: This type of channel has been widely considered in wireless communication studies. The ill condition arises from the symmetric MIMO size.
  • Symmetric i.i.d. Rician: The ill condition arises not only from the symmetric MIMO size but also from the strong LoS correlation.
@@
Figure 2: PDF of $(\theta_{N-1}+\theta_{N-2})/(-\theta_{0}+\theta_{N-2})$ for the PIA approach in both ELAA channels and stationary channels.
@@ V-C1 Experiment 1
This experiment aims to investigate the performance of the proposed PIA/e-PIA approaches in comparison to the theoretical expectation in Theorem 1. Here we consider the ZF precoding, i.e., $\mathbf{A}=\mathbf{H}\mathbf{H}^{H}$. The performance using the RZF will be shown in Experiment 4$\sim$5.
@@
Fig. 2 demonstrates the PDF of (85), where the solid lines stand for the ELAA channels and the dash lines for the stationary channels. It is shown that $\theta_{N-1}$ almost always falls in the expected interval. Only in the LoS-dominated ELAA, $\theta_{N-1}$ is sometimes slightly larger than $-\theta_{N-2}$ (but still negative). This is because the channel of LoS-dominated ELAA can be healthy-conditioned with a small probability. In this case, the assumptions in Theorem 3 (i.e., $\lambda_{0}\gg\lambda_{N-2}$ and $\lambda_{1}\gg\lambda_{N-2}$) no longer holds. Fortunately, Fig. 2 shows that $\theta_{N-1}$ is still close to the target interval. Therefore, we can say that the PIA approach successfully fulfills Theorem 1, such that $\kappa(\mathbf{R})=(\theta_{0})/(\theta_{N-2})$.
Figure 3: PDF of $(\theta_{0})/(\lambda_{1})$ for the PIA approach in LoS-dominated ELAA when $M=64$, $N=16$ and i.i.d. Rician channel when $M=N=128$.
Then, we investigate the behavior of $\theta_{0}$ and $\theta_{N-2}$ compared to the optimal performance in (17).
We first study the behavior of $\theta_{0}$, where the PIA alone is expected to successfully reduce $\theta_{0}$ to $\lambda_{0}$. As shown in Fig. 3, the red line stands for the PDF of $(\theta_{0})/(\lambda_{1})$, while the black line stands for $(\lambda_{0})/(\lambda_{1})$. We only show the case where $\lambda_{0}\gg\lambda_{1}$: LoS-ELAA and i.i.d. Rician channel. It is demonstrated that in both cases, $(\theta_{0})/(\lambda_{1})$ is less than $1$ dB, providing more than $10$ dB improvement in average compared to $(\lambda_{0})/(\lambda_{1})$. This shows the PIA approach successfully achieves the theoretical expectation.
Secondly, we examine $\theta_{N-2}$ compared to $\lambda_{N-2}$. Fig. 4 shows both the PIA/e-PIA as the list size $L$ increases from $1$ to $N$. We consider three cases where $\lambda_{N-1}\ll\lambda_{N-2}$: mixed LoS/non-LoS ELAA, i.i.d. Rayleigh and i.i.d. Rician channel. Here we consider $M=N=128$. Later in Experiment 3, we will show the case when $M=N=512$.
As shown in Fig. 4, we have $(\lambda_{N-1})/(\lambda_{N-2})$ averagely stays around $-10$ dB and can reach $-20$ dB. This shows that $\lambda_{N-1}$ compared to $\lambda_{N-2}$ alone is significantly increasing the condition number of $\mathbf{A}$. This is due to the symmetric MIMO size. In all three cases, the PIA approach (i.e., $L=1$) brings significant improvement: $(\theta_{N-2})/(\lambda_{N-2})$ is around $-4$ dB in average with a left tail of $-15$ dB. $(\theta_{N-2})/(\lambda_{N-2})$ is not so close to 00 dB as in Fig. 3. This is because $\hat{\mathbf{u}}_{N-1}$ is not well estimated, as expected in Section IV-B.
As the list size increases to $L=4$ or $8$, $(\theta_{N-2})/(\lambda_{N-2})$ is around $-2$ dB in average with a left tail of only $-8$ dB. Moreover, the e-PIA approach with $L=8$ already achieves close performance to $L=N$. This means a reasonable list size of e-PIA is enough to improve the performance beyond the PIA. Overall, the e-PIA demonstrates significant improvement with a reasonable list size. Later in Experiment 3$\sim$5, the list size of the e-PIA approach is set as $L=8$.
@@ V-C2 Experiment 2
Figure 4: PDF of $(\theta_{N-2})/(\lambda_{N-2})$ of the PIA/e-PIA approaches w.r.t. the list size $L$ when $M=N=128$ in (a) mixed LoS/non-LoS ELAA ($i=46$); (b) i.i.d. Rayleigh channel ($i=34$); (c) i.i.d. Rician channel ($i=34$).
This experiment aims to investigate the scalability of the SR-$1$R by showing the computation latency, as shown in Fig. 5. The computation device is Nvidia RTX A6000. The results are averaged based on $5,000$ implementations. Notably, substantial efforts have been made to accelerate matrix inversion on GPU platforms (e.g., [45, 46]). Here, the matrix partitioning is employed to accelerate the computation of GS and SSOR [46, Theorem 1]. It is shown that the PIA roughly increases from around $0.7$ ms to $1.3$ ms as $N$ increases from $16$ to $512$. The computation latency of the e-PIA is higher than the PIA, e.g., $3.8$ ms when $N=512$, yet still at the same level. This coincides with our scalability analysis. In conclusion, both the PIA/e-PIA show outstanding scalability. In comparison, the GS and SSOR reach $25$ ms and $31$ ms when $N=512$, respectively.
Fig. 5 also includes the latency of one iteration for the Schulz method and the extra latency to calculate the linear precoder, as shown by the black dash lines. They will be used to compared the overall latency for iterative linear precoding later in Experiment 5.
Figure 6: CDF of condition number for the PIA/e-PIA approaches compared to the preconditioning baselines in i.i.d. Rayleigh channel and i.i.d. Rician channel when $M=N=128$.
@@ V-C3 Experiment 3
This experiment aims to investigates the condition number improvement of the PIA/e-PIA approaches, as shown in Fig 6$\sim$8. Specifically, the black solid line stands for the cumulative distribution function (CDF) of $\kappa(\mathbf{A})$ (denoted by ‘Original’), while the black dash line stands for the CDF of $(\lambda_{1})/(\lambda_{N-2})$ as the performance bound of the SR-$1$R (denoted by ‘Bound’). The red lines with circle markers stand for the proposed PIA/e-PIA approaches, while the blue lines stand for the preconditioning baselines.
We first discuss the case where the channel is stationary. Fig. 6(a) and Fig. 7(a) shows the CDF of the condition number in i.i.d. Rayleigh channel when $M=N=128$. It is shown that the PIA significantly reduces the tail of the CDF. For example, when the CDF function is $0.9$, the PIA is around $5$ dB smaller than the GS and SSOR preconditioning. In communication systems, this indicates that the GS and SSOR will need more iterations to handle the extremely large condition number such that the communication error can be minimized in average. In addition, the e-PIA extends this advantage to around $10$ dB. As for the Jacobi preconditioning, it hardly changes the condition number from the original case.
When the channel is i.i.d. Rician, the improvement brought by the PIA/e-PIA is even more significant. As shown in Fig. 6(b) and Fig. 7(b), the PIA/e-PIA reduces the condition number for around $25$ dB. Moreover, in both types of channels, the e-PIA is close to the performance bound.
Fig. 8 demonstrates the CDF of the condition number when the channel is spatially non-stationary. When the channel is LoS-dominated, the PIA can reach the performance bound already, as shown in Fig. 8(a). In this case, the PIA improves the condition number for around $6$ dB compared to the GS preconditioning. When the channel is mixed LoS/non-LoS, the PIA/e-PIA shows similar behavior to the case of i.i.d. Rayleigh fading, i.e., the PIA and e-PIA achieve around $5$ dB and $10$ dB improvement when the CDF function is $0.9$, respectively.
Figure 7: CDF of condition number for the PIA/e-PIA approaches compared to the preconditioning baselines in i.i.d. Rayleigh channel and i.i.d. Rician channel when $M=N=512$.
Figure 8: CDF of condition number for the PIA/e-PIA approaches compared to the preconditioning baselines in LoS-dominated ELAA channel ($M=64$, $N=16$) and mixed LoS/non-LoS ELAA channel ($M=N=128$).
V-C4 Experiment 4
This experiment aims to investigate the behavior of iterative linear precoding with the RZF compared to the ZF. To facilitate the analysis, define the SNR as the ratio of the transmission power $\mathcal{E}$ over the noise power $\sigma_{v}^{2}$, i.e., $(\mathcal{E})/(\sigma_{v}^{2})$. This is also referred to as the normalized transmission power.
@@ V-C4 Experiment 4
Overall, the RZF improves the SER, but does not change the behavior of iterative linear precoding In Experiment 5, the RZF will be employed to for further investigation.
Figure 9: Average SER of the e-PIA ($i=46$), the Schulz method ($i=50$) and the preconditioning baselines ($i=50$) for the RZF and ZF w.r.t. $(\mathcal{E})/(\sigma_{v}^{2})$ when $M=N=128$ in mixed LoS/non-LoS ELAA.
V-C5 Experiment 5

### v4

@@ III-B Optimality Study
This result indicates the best matrix condition that can be possibly achieved via the SR-11R approach.
@@
IV Power Iteration-Assisted Approach for The Implementation of SR-11R
@@ IV-A The Proposed PIA Approach
Algorithm 1 Pseudo code of the PIA for the SR-11R
@@
6: Form the matrix $\mathbf{R}$ via (1) and compute $\mathbf{R}^{-1}$ through the Schulz method specified in (9);
@@
Finally, the proposed PIA approach for SR-11R implementation is summarized in Algorithm 1.
@@ IV-B Convergence of The PIA Approach
7: Let $\lambda_{\textsc{old}}=\lambda$ and goto step 33;
@@ IV-C Enhanced PIA Approach
Define $\mbox{$\mathcal{\phi}$}(i)\triangleq(\phi_{1}^{i},\phi_{2}^{i},...,\phi_{N-1}^{i})^{T}/\phi_{0}^{i}$ and represent the term $\widetilde{\lambda}(i)$ in (73) as the following vector form
$\widetilde{\lambda}(i)=\mbox{$\mathcal{\phi}$}^{T}(i)\mathbf{c}^{\prime}.$ (75)
@@
$l^{\star}=\underset{l}{\arg\min}|\mbox{$\mathcal{\phi}$}^{T}(i)\mathbf{c}^{\prime}_{l}|,~_{l=0,...,L-1}.$ (76)
@@ IV-D Scalability Analysis for SR-1R
In this subsection, we analyze the scalability of iterative linear precoding with the SR-11R in terms of both the overall complexity and the algorithm-depth. Then, we compare the SR-11R to the preconditioning baselines. The complexity analysis reflects the computational cost of the SR-11R; while the algorithm-depth reflects the minimum computing steps (i.e., the shortest processing time) when employing parallel computing [34]. This measure is particularly relevant because compatibility with parallel computing is a key advantage of iterative matrix inversion methods [1]. It will be shown that both the PIA/e-PIA algorithms exhibit favorable compatibility with parallel computing.
@@
  • Step $6$. This step belongs to the iterative matrix inversion. Each iteration costs two matrix-matrix multiplications and a matrix-matrix deduction. When the iteration nmber is $i$, this yields the complexity of $(2N^{3}+N^{2})i$ and the algorithm-depth of $2i\log_{2}N$.
@@
Summing up these steps apart from Step $6$ yields the overall algorithm depth of the SR-11R:
$(\tau_{0}+\tau_{N-1}+7)\log_{2}N+\log_{2}M.$ (79)
Step $6$ is excluded in (79), as it belongs to the iterative matrix inversion itself rather than to the preconditioning or regularization procedures. A good regularization/preconditioning technique accelerates this process. In Section V-C, we will demonstrate the iterations needed for convergence and computational latency when the iterative matrix inversion converges (see Table I).
The overall computational complexity of the SR-11R is:
$N^{3}+(\tau_{0}+\tau_{N-1}+M+7)N^{2}+(M+\log_{2}N+4)N.$ (80)
We observe that this complexity level is at the same level as that of direct inversion, which is $\mathcal{O}(N^{3})$ [26]. However, as emphasized in the introduction, this additional complexity is a common trade-off to enhance compatibility with parallel computing architectures [35, 36].
When employing the e-PIA, Algorithm 1 is performed in parallel for $L$ times. Therefore, the complexity level of the e-PIA is $\mathcal{O}(LN^{3})$; while the algorithm-depth level remains the same as the PIA.
The preconditioning baselines usually involves the inversion of a triangular matrix, i.e., a complexity of $\mathcal{O}(N^{2})$ [37]. This is done by the Gauss-Jordan elimination, and constitutes $N$ iterations [38]. Hence, the algorithm-depth is $\mathcal{O}(N)$. After taking the necessary extra steps (e.g., (78)) into consideration, the overall complexity of preconditioning baselines is dominated by the Gershgorin circle theorem, i.e., $\mathcal{O}(N^{3})$. While the algorithm depth is dominated by the triangular matrix inversion, i.e., $\mathcal{O}(N)$. Hence, the SR-11R reduces the algorithm depth from linear level to logarithm level of $N$, while maintaining the complexity at the same level as the preconditioning baselines.
In Section V, we will demonstrate the scalability of the SR-11R in terms of the computation latency when parallel computing is supported (see Fig. 3) as well as the overall latency of iterative linear precoding when RZF performance is achieved (see Table I).
@@ V Simulation Results and Discussion
In this section, computer simulations are carried out to demonstrate the performance of both the PIA and e-PIA approaches for the SR-11R. Specifically, Section V-A shows the implementation details of the SR-11R; Section V-B details the baselines and simulation setup; Section V-C shows simulation results.
Figure 2: Results for Experiment 1: (a) CDF of $(\theta_{N-1}+\theta_{N-2})/(-\theta_{0}+\theta_{N-2})$ for the PIA approach; CDF of condition number compared to the preconditioning baselines in (b) i.i.d. Rayleigh channel, (c) i.i.d. Rician channel, (d) LoS-dominated ELAA, (e) mixed LoS/non-LoS ELAA.
V-A Implementation of the SR-11R
@@
(81) means the increment approaches zero in two cases: when $i\to\infty$, or when $(\lambda_{n})/(\lambda_{0})\approx 1$. When $\mathbf{u}$ is badly initialized, i.e., $\lambda_{n}\approx\lambda_{0}$, it is hard to determine the convergence. This brings troubles to the PIA and e-PIA approaches to calculate $\hat{\mathbf{u}}_{N-1}$ .
@@ V-B Baselines and Simulation Setup
We consider two types of baselines: 1) iterative matrix inversion, where we use the Schulz method (see (9)) [2]; 2) preconditioning, where we use the Jacobi preconditioning (see (12)) [15, Eq. (10)], Gauss-Seidel (GS) preconditioning (see (13)) [23, Eq. (14)], and Symmetric successive over-relaxation (SSOR) (see (14)).
We consider two types of stationary ill-conditioned channel and two types of ELAA channel in the simulation. For stationary channels, we consider 1) Symmetric i.i.d. Rayleigh, where the ill condition arises from the symmetric MIMO size; 2) Symmetric i.i.d. Rayleigh, where the ill condition also arises from the strong LoS correlation. In most cases, we set $M=N=512$, as iterative matrix inversion is particularly beneficial in large-scale MIMO systems.
For ELAA channels, the channel model in [11] is adopted for the ELAA channel in appreciation of its comprehensive consideration of mixed LoS/non-LoS states, shadowing effects and path-loss. We consider the height of users and the base station to be $1.5$ m and $10$ m, respectively. Specifically, the considered ELAA channels include: 1) LoS-dominated ELAA, where the ill condition arises from the correlation of the LoS paths. We consider $M=64$, $N=16$ (single user), and the vertical distance from the user to the ELAA is $40$ m; 2) Symmetric Mixed LoS/non-LoS ELAA, where the ill condition arises from both the correlation of the LoS paths and the symmetric MIMO size. The MIMO size is configured as: $M=N=128$ (88 users with $16$ antennas each), and the vertical distance from users to the ELAA is $40$ m. The users are equally spaced along a parallel line to the base station antenna array with a maximal inter-user distance of $10$ m.
@@ V-C Simulation Results and Discussion
Figure 4: Average SER of the e-PIA ($i=46$) and the baselines ($i=50$) for the RZF and ZF w.r.t. $(\mathcal{E})/(\sigma_{v}^{2})$ in mixed LoS/non-LoS ELAA.
Figure 5: Average SER of the PIA/e-PIA approaches as the iteration number increases compared to exact RZF performance in (a) i.i.d. Rayleigh channel, (b) i.i.d. Rician channel, (c) LoS-dominated ELAA, (d) mixed LoS/non-LoS ELAA.
@@ V-C1 Experiment 1
[…]
@@
Fig. 3(a) demonstrates the cumulative distribution function (CDF) of (83), where the solid lines stand for the ELAA channels and the dash lines for the stationary channels. It is shown that the measurement in (83) falls in the expected interval. Hence, the PIA approach successfully fulfills Theorem 1, such that $\kappa(\mathbf{R})=(\theta_{0})/(\theta_{N-2})$.
[…]
We first discuss the case where the channel is stationary. Fig. 3(b) shows the CDF of the condition number in i.i.d. Rayleigh channel when $M=N=512$. It is shown that the PIA significantly reduces the tail of the CDF. For example, when the CDF function is $0.9$, the PIA is around $5$ dB smaller than the GS and SSOR preconditioning. In communication systems, this indicates that the GS and SSOR will need more iterations to handle the extremely large condition number such that the communication error can be minimized in average. In addition, the e-PIA extends this advantage to around $10$ dB. As for the Jacobi preconditioning, it hardly changes the condition number from the original case.
When the channel is i.i.d. Rician, the improvement brought by the PIA/e-PIA is even more significant. As shown in Fig. 3(c), the PIA/e-PIA reduces the condition number for around $25$ dB. Moreover, in both types of channels, the e-PIA is close to the performance bound.
[…]
@@ V-C2 Experiment 2
This experiment aims to investigate the scalability of the SR-11R by showing the computation latency, as shown in Fig. 3. The computation device is Nvidia RTX A6000. The results are averaged based on $5,000$ implementations. Notably, substantial efforts have been made to accelerate matrix inversion on GPU platforms (e.g., [39, 40]). Here, the matrix partitioning is employed to accelerate the computation of GS and SSOR [40, Theorem 1]. It is shown that the PIA roughly increases from around $0.7$ ms to $1.3$ ms as $N$ increases from $16$ to $512$. The computation latency of the e-PIA is higher than the PIA, e.g., $3.8$ ms when $N=512$, yet still at the same level. This coincides with our scalability analysis. In conclusion, both the PIA/e-PIA show outstanding scalability. In comparison, the GS and SSOR reach $25$ ms and $31$ ms when $N=512$, respectively.
Fig. 3 also includes the latency of one iteration for the Schulz method and the extra latency to calculate the linear precoder, as shown by the black dash lines. They will be used to compared the overall latency for iterative linear precoding later in Experiment 4.
@@ V-C3 Experiment 3
This experiment aims to investigate the behavior of iterative linear precoding with the RZF compared to the ZF. We choose the case of mixed LoS/non-LoS ELAA as representative. In other cases, the phenomenon observed is similar. To facilitate the analysis, define the SNR as the ratio of the transmission power $\mathcal{E}$ over the noise power $\sigma_{v}^{2}$, i.e., $(\mathcal{E})/(\sigma_{v}^{2})$. This is also referred to as the normalized transmission power.
@@
Overall, the RZF improves the SER, but does not change the behavior of iterative linear precoding In Experiment 4, the RZF will be employed to for further investigation.
V-C4 Experiment 4